Give an example of 2 × 2 matrix with non-zero entries that has no inverse.


1
Expert's answer
2021-05-26T14:17:44-0400

If the determinant of a square matrix A is 0, then A has no inverse.

For example


"A=\begin{pmatrix}
 2 & 1 \\
 -2 & -1
\end{pmatrix}"

"\text{det}A=\begin{vmatrix}
 2 & 1 \\
 -2 & -1
\end{vmatrix}=2(-1)-1(-2)=0"

The 2x2 matrix A has no inverse.